A man invested #20,000 in bank A and #25,000 in bank B at the beginning of a year. Bank A pays simple interest at a rate of y% per annum and B pays 1.5y% per annum. If his total interest at the end of the year from the two banks was #4,600 . Find the value of y.


1
Expert's answer
2021-09-07T10:08:44-0400

Interest from bank A : "\dfrac{y}{100}\times 20000=200y"


Interest from bank A : "\dfrac{1.5y}{100}\times 25000=375y"

His total interest at the end of the year from the two banks


"200y+375y=4600"

Solve for "y"


"575y=4600"

"y=\dfrac{4600}{575}"

"y=8"